# Testing inflation on all scales: a case study with $\alpha$-attractors

### v1

@@ 1.1 Methodology
For each point of the parameter space, we use the publicly available PyTransport package [30, 31, 32] to numerically compute the primordial 2- and 3-point correlation functions, see Eqs. (2.4) and (3.1). Note that enhanced fluctuations on small scales require deviations from single-field slow-roll [33], and therefore it is mandatory to compute the inflationary power spectrum numerically. Our methodology is then articulated into four parts:
@@
  • Theoretical consistency checks. Recently, there has been a lot of debate on the importance of 1-loop corrections in inflationary models leading to enhanced fluctuations on small scales [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. Available numerical codes operate at tree-level. If loop corrections are not important, it is safe to use their output to correctly evaluate the small-scale phenomenology. If not [19, 62], more work needs to be done to obtain correct inflationary predictions [^1] For recent developments with lattice simulations see Refs. [63, 64].. As a first step towards a systematic check of whether perturbativity holds at peak scales, we estimate the size of 1-loop corrections to the 2-point correlation function due to cubic interactions, and compare it with the tree-level result. We also implement a second theoretical test on the parameter space, by ensuring that PBHs are not overproduced.
@@ 1.2 A case study with hybrid $\alpha$-attractors
where a linear term (not present in standard hybrid-inflation models) has been introduced. This ensures that topological defects, e.g. domain walls, are not produced, while also addressing the problem of eternal inflation [21]. The dynamics and phenomenology of the model (1.2) have been extensively studied in Ref. [21], with emphasis on the possibility of enhancing the scalar power spectrum on small scales, and hence produce PBHs and large second-order GWs.
@@ 2.1 A brief review of background and perturbations dynamics
The introduction of a linear term in Eq. (1.2) ensures that topological defects, e.g. domain walls, are not produced, while also addressing the problem of eternal inflation [21]. Without loss of generality [^4] Note that we could equally assume that $d>0$, in which case $\chi_{\text{valley}}<0$ and, for $\phi<\phi_{\text{c}}$, the inflaton rolls towards the minimum at $\chi_{\text{min}}=-\chi_{0}-d/2+\mathcal{O}(d^{2})$. Also in this case, one has $V(\phi_{\text{min}},\chi_{\text{min}})=-M^{2}d\chi_{0}+\mathcal{O}(d^{2})<0$, which can be offset by introducing a constant, positive shift in the potential (1.2), see footnote 5. we assume $d<0$, in which case $\chi_{\text{valley}}$ is a small, positive number. When $\phi<\phi_{\text{c}}$, the $\chi$ field becomes tachyonic, i.e. its squared mass (2.1) turns negative, and evolves away from $\chi_{\text{valley}}$, towards the global potential minimum at $\phi=0$ and $\chi=\chi_{\text{min}}$. For $\chi_{0}\ll 1$, $\chi$ is very heavy ($m_{\chi}^{2}\gg H^{2}$) and inflation ends abruptly. On the other hand, for larger values of $\chi_{0}$, $\chi$ is allowed to slow roll and drive a second phase of inflation. This is the regime Ref. [21] investigates, and which will also be our focus.
@@
where higher-order coefficents can be found by solving $V_{\chi}|_{\chi_{\text{min}}}=0$ order by order555 At the minimum, the potential is given by the negative, small value $V(\phi_{\text{min}},\chi_{\text{min}})=M^{2}d\chi_{0}+\mathcal{O}(d^{2})$. In order to have vanishing potential energy at the minimum, we consider to add a constant, positive contribution to the potential (1.2), $\Delta V=M^{2}|d|\chi_{0}$. Since this term would have no consequences for the inflationary dynamics, we have not explicitly included it in Eq. (1.2)..
@@ 2.2 Calibrating the horizon-crossing of scales
The results are shown in Fig. 2, where one sees that there is a small variation of $\Delta N_{\text{CMB, inst rh}}$ over the parameter space, $54.597\leq\Delta N_{\text{CMB, inst rh}}\leq 55.901$, with the smallest value corresponding to the model $\{\chi_{0}=2.6,\,d=-10^{-6}\}$ and the largest to $\{\chi_{0}=2.474,\,d=-10^{-6}\}$. The rapid variation of $\Delta N_{\text{CMB, inst rh}}$ in the top-right corner of the parameter space is due to the fact that for those models multi-field effects are already relevant at CMB scales and the scale of inflation is lower as a consequence (see footnote 7).
Starting from the results for $\Delta N_{\text{CMB, inst rh}}$, we find the allowed values of $\Delta N_{\text{CMB}}$ for each model in the parameter space by including the effect of reheating as in Eq. (2.5). In particular, for each model $\Delta N_{\text{CMB, min}}\leq\Delta N_{\text{CMB}}\leq\Delta N_{\text{CMB, inst rh}}$, where $\Delta N_{\text{CMB, min}}$ corresponds to the case reheating lasted $\Delta\tilde{N}_{\text{rh, max}}$ e-folds [^8] We numerically compute the maximum duration of reheating using Eq. (2.6). We obtain $37.86\leq\Delta\tilde{N}_{\text{rh, max}}\leq 39.62$ over the whole parameter space, with smaller values in the top-right corner. This is again a consequence of having smaller $H$ in those models (see footnote 7), which in turns lowers $\rho_{\text{end}}$.. From the allowed range of $\Delta N_{\text{CMB}}$ computed for each point in the parameter space, we select three values which are compatible with all models and which will be the focus of our analysis for this work: $\Delta N_{\text{CMB}}=\{50,52.5,54\}$. The three selected values correspond to different duration of reheating over the parameter space, which for each model is given by $\Delta\tilde{N}_{\text{rh}}=4\left(\Delta N_{\text{CMB, inst rh}}-\Delta N_{\text{CMB}}\right)$, see Eq. (2.5). Specifically,
@@ 2.3 CMB anisotropies constraints
Figure 3: Large-scale constraints on the $(\chi_{0},\,d)$ parameter space for different values of $\Delta N_{\text{CMB}}$, with fixed $\{\alpha=1,\,\tilde{m}=0.3,\,\tilde{g}=0.8\}$. The white areas are those consistent with both CMB and $\mu$-constraints at least at $95\%$ C.L., see Sec. 2.3 and 2.4 for more details. The red contours identify models whose scalar power spectrum peaks at LISA and PTA scales, $k_{\text{LISA}}=3\times 10^{11}\,\text{Mpc}^{-1}$ and $k_{\text{PTA}}=10^{7}\,\text{Mpc}^{-1}$.
In Fig. 3 we represent the parameter space constraints stemming from the measurements above. In particular, the white regions are those complying with both constraints at least at $95\%$ C.L.. The red lines mark models whose peak in $\mathcal{P}_{\zeta}(k)$ is at LISA and PTA scales, see Sec. 4.2. They are not exactly vertical in the $(\chi_{0},\,d)$ plane, indicating that the peak position, while being strongly correlated with $\chi_{0}$, also mildly depends on $d$. By observing how the white area changes with $\Delta N_{\text{CMB}}$, we conclude that, assuming $w=0$, an extended reheating stage (i.e. smaller $\Delta N_{\text{CMB}}$) shrinks the available region of parameter space. This is because lowering $\Delta N_{\text{CMB}}$ moves the peak in $\mathcal{P}_{\zeta}(k)$ closer to $k_{\text{CMB}}$, see how the red contours change from one panel to the other.
@@ 3 Non-Gaussianity and perturbativity
  • a squeezed, with $k_{1}=k_{t}/20,\,k_{2}=k_{t}/2,\,k_{3}=9k_{t}/20$ (or $\alpha=0.1$ and $\beta=0.9$).
@@
In Fig. 4 the results are represented as a function of $k_{t}$, with $k_{t}$ ranging from the CMB pivot scale to the one that crossed the horizon at the end of inflation. First, we note that on large scales the reduced bispectrum is small, e.g. ${f_{\text{NL,\,sq}}}(k_{t}=k_{\text{CMB}})=0.01768$, well within Planck constraints [39], regardless of the configuration. At the pivot scale, the variation in the value of $f_{\text{NL}}$ between the four configurations considered is less than $4\%$. The value of $f_{\mathrm{NL}}$ in the squeezed configuration is in agreement with the prediction from Maldacena’s consistency relation [84], $f_{\text{NL, CR}}=(5/12)(1-n_{\text{s}})=0.01762$. This is due to the fact that, when the large scales crossed the horizon, the dynamics of the model is essentially SFSR, driven by the slowly-rolling field $\phi$. While the results represented in Fig. 4 are for a single model, we have computed the (equilateral and squeezed) reduced bispectrum on large scales over the $(\chi_{0},\,d)$ parameter space compatible with the large-scale constraints discussed in Sec. 2. We found $f_{\text{NL,\,eq}}(k_{t}=3k_{\text{CMB}})=\left[0.012,0.019\right]$ and $f_{\text{NL,\,sq}}(k_{t}=3k_{\text{CMB}})=\left[0.009,0.019\right]$, compatible with CMB constraints [39]. Hence, the computation of the bispectrum on large scales does not allow us to further constrain the parameter space. It is important to point out that this study of non-Gaussianity for models compatible with large-scale constraints does not depend on the duration of reheating, i.e. the specific chosen value of $\Delta N_{\text{CMB}}$. All the models in the $(\chi_{0},d)$ parameter space region compatible with the large-scale constraints on $n_{\text{s}}$, $\alpha_{s}$ and $\mu$-distortion (see Fig. 3), will have negligible non-Gaussianity on CMB scales regardless of $\Delta N_{\text{CMB}}$. This is because models complying with large-scale constraints effectively behave as in SFSR inflation at those scales.
Moving towards smaller scales, one observes in Fig. 4 that the reduced bispectrum is constant for modes that correspond to the peak in the scalar power spectrum. This applies regardless of the triangle configuration, and the $f_{\mathrm{NL}}$ plateau is configuration-independent. The small amplitude of $f_{\mathrm{NL}}$, e.g. $f_{\text{NL}}(k_{t}=k_{\text{peak}})=0.45883$, suggests that non-linearities due to cubic interactions are subdominant and do not pose a threat to perturbativity for this model, see Eq. (3.6) and the related discussion.
@@ 4.1 Primordial black hole formation
We represent our results in Fig. 7, where each panel corresponds to a different choice for $\Delta N_{\text{CMB}}$. Our results show that imposing $f_{\text{PBH}}\leq 1$ yields significantly stronger constraints on the parameter space with respect to the approximate criterion which is often used, $\mathcal{P}_{\zeta}(k_{\text{peak}})\lesssim 0.01$.
@@ 4.2 Scalar-induced gravitational waves
We assume the SIGW are produced at horizon re-entry during radiation domination, and the fraction of SIGW energy density is [6, 7]
@@ 5 Discussion
Let us first comment on the methodology. In this work, we have illustrated the effect of a matter-dominated reheating stage by considering three different values of $\Delta N_{\text{CMB}}$. These have been carefully selected by ensuring that they are compatible with the parameter space we consider, see Sec. 2.2. In future work, we plan to describe reheating following the parametrisation introduced in Ref. [132], and systematically include the reheating parameter alongside other inflationary parameters. Furthermore, while we have constrained the parameter space by means of marginalised measurements/upper bounds obtained from current data, Bayesian inference for inflation can also be performed (see e.g. Refs. [133, 29]). This would allow us to derive the current posterior distribution for the inflationary parameters, and to perform Bayesian model comparison. Finally, as commented at the end of Sec. 4.1, for parameter-space regions leading to peaks on very small scales, e.g. $k_{\text{peak}}\gtrsim 10^{16}\,\text{Mpc}^{-1}$, the inclusion of observational limits on the PBH abundance as a function of the PBH mass would potentially yield to further constraints.

### v2

@@ 1.1 Methodology
For each point of the parameter space, we use the publicly available PyTransport package [30, 31] to numerically compute the primordial 2- and 3-point correlation functions, see Eqs. (2.4) and (3.1). PyTransport provides a numerical implementation of the transport approach to primordial correlators, see e.g. Ref. [32]. Note that enhanced fluctuations on small scales require deviations from single-field slow-roll [33], and therefore it is mandatory to compute the inflationary power spectrum numerically. Our methodology is then articulated into four parts:
@@
  • Theoretical consistency checks. Recently, there has been a lot of debate on the importance of 1-loop corrections in inflationary models leading to enhanced fluctuations on small scales [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Available numerical codes operate at tree-level. If loop corrections are not important, it is safe to use their output to correctly evaluate the small-scale phenomenology. If not [19, 63], more work needs to be done to obtain correct inflationary predictions [^1] For recent developments with lattice simulations see Refs. [64, 65].. As a first step towards a systematic check of whether perturbativity holds at peak scales, we estimate the size of 1-loop corrections to the 2-point correlation function due to cubic interactions, and compare it with the tree-level result. We also implement a second theoretical test on the parameter space, by ensuring that PBHs are not overproduced.
@@ 1.2 A case study with hybrid $\alpha$-attractors
where a linear term (not present in standard hybrid-inflation models) has been introduced. This ensures that topological defects, e.g. domain walls, are not produced, while also addressing the problem of eternal inflation [21]. The dynamics and phenomenology of the model (1.2) have been extensively studied in Ref. [21], with emphasis on the possibility of enhancing the scalar power spectrum on small scales, and hence produce PBHs and large second-order GWs [^3] For scenarios where the fields of the hybrid $\alpha$-attractor model (1.2) are coupled with a dark fermion see Refs. [71, 72]..
@@ 2.1 A brief review of background and perturbations dynamics
The introduction of a linear term in Eq. (1.2) ensures that topological defects, e.g. domain walls, are not produced, while also addressing the problem of eternal inflation [21]. Without loss of generality [^5] Note that we could equally assume that $d>0$, in which case $\chi_{\text{valley}}<0$ and, for $\phi<\phi_{\text{c}}$, the inflaton rolls towards the minimum at $\chi_{\text{min}}=-\chi_{0}-d/2+\mathcal{O}(d^{2})$. Also in this case, one has $V(\phi_{\text{min}},\chi_{\text{min}})=-M^{2}d\chi_{0}+\mathcal{O}(d^{2})<0$, which can be offset by introducing a constant, positive shift in the potential (1.2), see footnote 6. we assume $d<0$, in which case $\chi_{\text{valley}}$ is a small, positive number. When $\phi<\phi_{\text{c}}$, the $\chi$ field becomes tachyonic, i.e. its squared mass (2.1) turns negative, and evolves away from $\chi_{\text{valley}}$, towards the global potential minimum at $\phi=0$ and $\chi=\chi_{\text{min}}$. For $\chi_{0}\ll 1$, $\chi$ is very heavy ($m_{\chi}^{2}\gg H^{2}$) and inflation ends abruptly. On the other hand, for larger values of $\chi_{0}$, $\chi$ is allowed to slow roll and drive a second phase of inflation. This is the regime Ref. [21] investigates, and which will also be our focus.
@@
where higher-order coefficents can be found by solving $V_{\chi}|_{\chi_{\text{min}}}=0$ order by order666 At the minimum, the potential is given by the negative, small value $V(\phi_{\text{min}},\chi_{\text{min}})=M^{2}d\chi_{0}+\mathcal{O}(d^{2})$. In order to have vanishing potential energy at the minimum, we consider to add a constant, positive contribution to the potential (1.2), $\Delta V=M^{2}|d|\chi_{0}$. Since this term would have no consequences for the inflationary dynamics, we have not explicitly included it in Eq. (1.2)..
@@ 2.2 Calibrating the horizon-crossing of scales
The results are shown in Fig. 2, where one sees that there is a small variation of $\Delta N_{\text{CMB, inst rh}}$ over the parameter space, $54.597\leq\Delta N_{\text{CMB, inst rh}}\leq 55.901$, with the smallest value corresponding to the model $\{\chi_{0}=2.6,\,d=-10^{-6}\}$ and the largest to $\{\chi_{0}=2.474,\,d=-10^{-6}\}$. The rapid variation of $\Delta N_{\text{CMB, inst rh}}$ in the top-right corner of the parameter space is due to the fact that for those models multi-field effects are already relevant at CMB scales and the scale of inflation is lower as a consequence (see footnote 8).
Starting from the results for $\Delta N_{\text{CMB, inst rh}}$, we find the allowed values of $\Delta N_{\text{CMB}}$ for each model in the parameter space by including the effect of reheating as in Eq. (2.5). In particular, for each model $\Delta N_{\text{CMB, min}}\leq\Delta N_{\text{CMB}}\leq\Delta N_{\text{CMB, inst rh}}$, where $\Delta N_{\text{CMB, min}}$ corresponds to the case reheating lasted $\Delta\tilde{N}_{\text{rh, max}}$ e-folds [^9] We numerically compute the maximum duration of reheating using Eq. (2.6). We obtain $37.86\leq\Delta\tilde{N}_{\text{rh, max}}\leq 39.62$ over the whole parameter space, with smaller values in the top-right corner. This is again a consequence of having smaller $H$ in those models (see footnote 8), which in turns lowers $\rho_{\text{end}}$.. From the allowed range of $\Delta N_{\text{CMB}}$ computed for each point in the parameter space, we select three values which are compatible with all models and which will be the focus of our analysis for this work: $\Delta N_{\text{CMB}}=\{50,52.5,54\}$. The three selected values correspond to different duration of reheating over the parameter space, which for each model is given by $\Delta\tilde{N}_{\text{rh}}=4\left(\Delta N_{\text{CMB, inst rh}}-\Delta N_{\text{CMB}}\right)$, see Eq. (2.5). Specifically,
@@ 2.3 CMB anisotropies constraints
Figure 3: Large-scale constraints on the $(\chi_{0},\,d)$ parameter space for different values of $\Delta N_{\text{CMB}}$, with fixed $\{\alpha=1,\,\tilde{m}=0.3,\,\tilde{g}=0.8\}$. The white areas are those consistent with both CMB and $\mu$-constraints at least at $95\%$ C.L., see Sec. 2.3 and 2.4 for more details. The red contours identify models whose scalar power spectrum peaks at the LISA and PTA pivot scale, $k_{\text{LISA}}=3\times 10^{11}\,\text{Mpc}^{-1}$ and $k_{\text{PTA}}=10^{7}\,\text{Mpc}^{-1}$ respectively
In Fig. 3 we represent the parameter space constraints stemming from the measurements above. In particular, the white regions are those complying with both constraints at least at $95\%$ C.L.. The red contours correspond to models whose $\mathcal{P}_{\zeta}(k)$ peaks exactly at the LISA and PTA pivot scales (see caption of Fig. 3), which are representative of the frequency bands the two experiments operate at. These lines guide the eye and show at a glimpse the approximate location of the power spectrum peak across the parameter space. They are not exactly vertical in the $(\chi_{0},\,d)$ plane, indicating that the peak position, while being strongly correlated with $\chi_{0}$, also mildly depends on $d$. By observing how the white area changes with $\Delta N_{\text{CMB}}$, we conclude that, assuming $w=0$, an extended reheating stage (i.e. smaller $\Delta N_{\text{CMB}}$) shrinks the available region of parameter space. This is because lowering $\Delta N_{\text{CMB}}$ moves the peak in $\mathcal{P}_{\zeta}(k)$ closer to $k_{\text{CMB}}$, see how the red contours change from one panel to the other.
@@ 3 Non-Gaussianity and perturbativity
  • a squeezed, with $k_{1}=k_{t}/20,\,k_{2}=k_{t}/2,\,k_{3}=9k_{t}/20$ (or $\alpha=0.1$ and $\beta=0.9$). Note that all three modes $k_{1}$, $k_{2}$ and $k_{3}$ are comparable, with the maximum squeezing given by $k_{1}/k_{2}=0.1$.
@@
In Fig. 4 the results are represented as a function of $k_{t}$, with $k_{t}$ ranging from the CMB pivot scale to the one that crossed the horizon at the end of inflation. First, we note that on large scales the reduced bispectrum is small, e.g. ${f_{\text{NL,\,sq}}}(k_{t}=k_{\text{CMB}})=0.01768$, well within Planck constraints [39], regardless of the configuration. At the pivot scale, the variation in the value of $f_{\text{NL}}$ between the four configurations considered is less than $4\%$. The value of $f_{\mathrm{NL}}$ in the squeezed configuration when $k_{t}$ is a long mode is in agreement with the prediction from Maldacena’s consistency relation [87], $f_{\text{NL, CR}}(k_{t}=k_{\text{CMB}})=(5/12)(1-n_{\text{s}})=0.01762$. This is due to the fact that, when the large scales crossed the horizon, the dynamics of the model is essentially SFSR, driven by the slowly-rolling field $\phi$, and the long mode $k_{1}\ll k_{2},k_{3}$ is adiabatic [93, 94]. While the results represented in Fig. 4 are for a single model, we have computed the (equilateral and squeezed) reduced bispectrum on large scales over the $(\chi_{0},\,d)$ parameter space compatible with the large-scale constraints discussed in Sec. 2. We found $f_{\text{NL,\,eq}}(k_{t}=3k_{\text{CMB}})=\left[0.012,0.019\right]$ and $f_{\text{NL,\,sq}}(k_{t}=3k_{\text{CMB}})=\left[0.009,0.019\right]$, compatible with CMB constraints [39]. Hence, the computation of the bispectrum on large scales does not allow us to further constrain the parameter space. It is important to point out that this study of non-Gaussianity for models compatible with large-scale constraints does not depend on the duration of reheating, i.e. the specific chosen value of $\Delta N_{\text{CMB}}$. All the models in the $(\chi_{0},d)$ parameter space region compatible with the large-scale constraints on $n_{\text{s}}$, $\alpha_{s}$ and $\mu$-distortion (see Fig. 3), will have negligible non-Gaussianity on CMB scales regardless of $\Delta N_{\text{CMB}}$. This is because models complying with large-scale constraints effectively behave as in SFSR inflation at those scales.
Moving towards smaller scales, one observes in Fig. 4 that the reduced bispectrum is constant for modes that correspond to the peak in the scalar power spectrum. This applies regardless of the triangle configuration, and the $f_{\mathrm{NL}}$ plateau is configuration-independent. The small amplitude of $f_{\mathrm{NL}}$, e.g. $f_{\text{NL}}(k_{t}=k_{\text{peak}})=0.45883$, suggests that non-linearities due to cubic interactions are subdominant and do not pose a threat to perturbativity for this model, see Eq. (3.6) and the related discussion. Note that Maldacena’s consistency relation would not reproduce the numerical squeezed bispectrum when $k_{t}$ is a short mode as the long mode of the configuration, $k_{1}\ll k_{2},k_{3}$, is not adiabatic.
@@ 4.1 Primordial black hole formation
We represent our results in Fig. 7, where each panel corresponds to a different choice for $\Delta N_{\text{CMB}}$. Our results show that imposing $f_{\text{PBH}}\leq 1$ yields significantly stronger constraints on the parameter space with respect to the approximate criterion which is often used, $\mathcal{P}_{\zeta}(k_{\text{peak}})\lesssim 0.01$. Indeed, the precise value of $f_{\text{PBH}}$ is model dependent [2], and a careful numerical computation is required in order to meaningfully constrain inflation.
@@ 4.2 Scalar-induced gravitational waves
Owing to our findings of Sec. 3, we assume that corrections to the spectrum of SIGW due to non-Gaussianity of $\zeta$ are sub-leading [127, 128, 129]. We assume the SIGW are produced at horizon re-entry during radiation domination, and the fraction of SIGW energy density is [6, 7]
@@ 5 Discussion
Let us first comment on the methodology. In this work, we have illustrated the effect of a matter-dominated reheating stage by considering three different values of $\Delta N_{\text{CMB}}$. These have been carefully selected by ensuring that they are compatible with the parameter space we consider, see Sec. 2.2. In future work, we plan to describe reheating following the parametrisation introduced in Ref. [140], and systematically include the reheating parameter alongside other inflationary parameters (see also Refs. [141, 142]). Furthermore, while we have constrained the parameter space by means of marginalised measurements/upper bounds obtained from current data, Bayesian inference for inflation can also be performed (see e.g. Refs. [143, 29]). This would allow us to derive the current posterior distribution for the inflationary parameters, and to perform Bayesian model comparison. Finally, as commented at the end of Sec. 4.1, for parameter-space regions leading to peaks on very small scales, e.g. $k_{\text{peak}}\gtrsim 10^{16}\,\text{Mpc}^{-1}$, the inclusion of observational limits on the PBH abundance as a function of the PBH mass would potentially yield to further constraints.